If a right triangle has sides of 3 and 5, what is the measurement of the other side?
Answer by edjones(8007)

1)
If 5 is the hypotenuse:
9+a^2=25
a^2=16
a=4
.
2)
If 5 is a leg:
9+25=34
c^2=34
c=sqrt(34) length of the hypotenuse.
.
Ed